with usual notations, show that
Cov aX( + bY cX, + dY) = acVar(X) + bdVar(Y) + ad( + bc)Cov( X, Y)

is it true or false? Give reasons.
1
Expert's answer
2020-03-10T12:01:35-0400

"Cov(aX+bY,cX+dY)=Cov(aX,cX+dY)+Cov(bY,cX +dY)="

"Cov(aX,cX)+Cov(aX,dY)+Cov(bY,cX)+Cov(bY,dY)="

"ac\ Cov(X,X)+ad \ Cov(X,Y)+bc\ Cov(Y,X)+bd \ Cov (Y,Y)="

"ac \ Var(X)+ad\ Cov(X,Y)+bc\ Cov(X,Y)+bd \ Var(Y)="

"=ac \ Var(X)+bd\ Var(Y)+(ad+bc)Cov(X,Y)"


So, "Cov(aX+bY,cX +dY)=ac \ Var(X)+bd \ Var(Y)+(ad+bc)Cov(X,Y)"


The following properties were used:

"Cov(X,Y)=Cov(Y,X)"

"Cov(aX,Y)=a\ Cov(X,Y)"

"Cov(X+Y,W)=Cov(X,W)+Cov(Y,W)"

"Cov (X,X)=Var(X)"


Answer: this statement is true.